# The Merging Galaxy Cluster Environment Affects the Morphology of Radio-AGN

### v1

@@ 6 Summary
In this work, we explored the effect of galaxy cluster mergers on the radio morphology of radio-AGN. Starting from a sample of 33 galaxy clusters identified as merging (8) or non-merging (25) by Bilton et al. (2020), we assembled a sample of 12 galaxy clusters (8 merging and 4 non-merging) that have LoTSS 144 MHz data. Our method to determine if the merging cluster environment has an effect on radio-AGN morphology was to have 5 experts classify the sources with a radio detection using Zooniverse, according to whether a source is weak compact (WC), strong compact (SC), compact extended (CE), extended (E), FR I or FR II, FR-morphology (FRM), bent tail (BT), or a head tail (HT). We combined the radio morphologies classes into C (WC+SC), CE, E, Jetted (FR I + FR II + FRM), and Disturbed (BT + HT), ( see §4.1), and we obtain the following results:
mergers do not exhibit a greater proportion of Compact, Compact Extended, and Jetted sources as compared to non-mergers (see §4.1)
non-mergers contain all the Extended classified sources in this study
mergers have a statistically significant (3) higher population proportion of Disturbed sources than non-mergers (see §4.1)
within merging clusters there is a marginally statistically significant (2) higher proportion of AGN that are identified only as radio-AGN compared to radio-AGN that are also identified as an optical AGN (see §4.2)
Overall we find that the merging environment seems to have an effect on the radio morphology of radio sources in that there are more disturbed (bent and head tails) sources in mergers. Our findings suggest the merging environment provides more opportunity for the jets to become disturbed.
@@
within merging clusters there is a marginally statistically significant (2) higher proportion of AGN that are identified only as radio-AGN compared to radio-AGN that are also identified as an optical AGN (see §4.2)
mergers do not exhibit greater proportions of radio-AGNs and radio-AGNs with an optical AGN detection compared to non-mergers.
This study presents significant opportunities for future work. One future direction is to follow a similar analysis as Bilton et al. (2020) and examine the spatial distribution of the classified sources in the phase space diagram. Of particular interest is the potential trends of radio classification at certain spatial cuts within the cluster. A similar future direction would be to investigate the phase space position of the AGNRad and AGNRad-Opt. An additional future direction is to extend this work to a larger sample of clusters and have a spectrum of cluster dynamical states. A last future direction would be the further investigation of the ”undetermined” or ”non-consensus” morphological classifications as there is a significant variation in the morphologies in these morphology classes.
@@ Acknowledgements
We would like to thank the anonymous referee for their useful comments that made this work better. We would also like to thank Drs. Abhijeet Borkar and Paul Martini for helpful discussions. This work was made possible by the support of The National Radio Astronomy Observatory and Green Bank Observatory which are facilities of the U.S. National Science Foundation operated under cooperative agreement by Associated Universities, Inc. M.R. acknowledeges support from the Deans’ Graduate Fellowship from the University of Notre Dame. Y.A.G. is supported by US National Science Foundation (NSF) grant AST 22-06053. This publication uses data generated via the Zooniverse.org platform, development of which is funded by generous support, including a Global Impact Award from Google, and by a grant from the Alfred P. Sloan Foundation.
Name | Cluster | Dstate | RA | DEC | Radioclass | RAGN | OAGN |
SDSS J031447.7+421321 | ABELL0426 | M | 48.698868 | 42.222702 | WC | False | False |
SDSS J114041.6+202034 | ABELL1367 | NM | 175.173660 | 20.342951 | WC | False | False |
SDSS J114224.4+200709 | ABELL1367 | NM | 175.602066 | 20.119303 | CE | False | False |
SDSS J114358.1+204822 | ABELL1367 | NM | 175.992096 | 20.806389 | WC | False | False |
SDSS J114358.9+200437 | ABELL1367 | NM | 175.995667 | 20.077028 | CE | False | False |
SDSS J114527.6+204825 | ABELL1367 | NM | 176.365280 | 20.807171 | CE | False | False |
SDSS J114548.8+203743 | ABELL1367 | NM | 176.453491 | 20.628616 | WC | False | False |
SDSS J114612.1+202329 | ABELL1367 | NM | 176.550766 | 20.391647 | SC | True | True |
SDSS J115148.4+202727 | ABELL1367 | NM | 177.951752 | 20.457710 | WC | False | False |
SDSS J125217.7+270507 | ABELL1656 | NM | 193.073837 | 27.085546 | WC | False | True |
SDSS J125351.5+285845 | ABELL1656 | NM | 193.464584 | 28.979395 | WC | False | False |
SDSS J125359.1+262638 | ABELL1656 | NM | 193.496307 | 26.444042 | WC | False | False |
SDSS J125455.1+272445 | ABELL1656 | NM | 193.729874 | 27.412704 | WC | False | False |
SDSS J125547.8+281521 | ABELL1656 | NM | 193.949280 | 28.256104 | WC | False | False |
SDSS J125651.1+265356 | ABELL1656 | NM | 194.213211 | 26.898890 | WC | False | True |
SDSS J125732.8+273637 | ABELL1656 | NM | 194.386826 | 27.610346 | WC | False | False |
SDSS J125835.1+273547 | ABELL1656 | NM | 194.646606 | 27.596390 | SC | False | False |
SDSS J125855.9+275000 | ABELL1656 | NM | 194.733185 | 27.833393 | CE | False | False |
SDSS J125939.1+285343 | ABELL1656 | NM | 194.912979 | 28.895521 | WC | False | False |
SDSS J130125.2+291849 | ABELL1656 | NM | 195.355286 | 29.313740 | CE | True | True |
Note. — Column headers are shorthand for the combined classification groups: Compact (WC+SC), Compact Extended, Extended, Jetted (FRI+FRII+FRM), and Disturbed (HT+BT). Rows describe the total radio-AGN and proportion of the classification within non-mergers and mergers using a beta-binomial Jeffreys Bayesian prior to determine the upper and lower bound uncertainty. corresponds to the posterior probability that a specific classification is greater in mergers compared to non-mergers, discussed in Section 3.6.
Note. — Columns are radio-AGN coinciding with an optical AGN determined in Bilton et al. (2020) (AGNRad-Opt) and only radio-AGN (AGNRad). Rows include the total number and the proportion with beta-binomial upper and lower bound uncertainties using a Jeffreys Bayesian prior across both dynamical states. corresponds to the posterior probability that a specific classification is greater in mergers compared to non-mergers, discussed in Section 3.6.
Note. — The first 20 rows of the full sample of consensus level 3 sources. The full table is provided with this article. Columns are the SDSS name of source Bilton et al. (2020) (Name), the name of the cluster that the source is in (Cluster), the dynamical state of the host cluster, Dstate, which is either non-merging (NM) or merging (M) as defined in Bilton et al. (2020), RA, Dec, the radio classification assigned through the Zooniverse classification done in this work (Radio class, see §3.2), whether or not the source is classified as a radio-AGN (RAGN, see §3.5), and whether or not the source is classified as an optical AGN in Bilton et al. (2020) (OAGN).
WC | SC | CE | E | FRI+II | FRM | BT | HT | Unc | |
TotNM | 79 | 19 | 17 | 7 | 1 | 2 | 1 | 0 | 0 |
TotM | 19 | 5 | 2 | 0 | 2 | 0 | 1 | 4 | 1 |
@@
C | CE | E | Jetted | Disturbed | |
TotNM | 16 | 5 | 2 | 2 | 0 |
TotM | 8 | 0 | 0 | 1 | 2 |
PNM | |||||
PM | |||||
0.69 | 0.04 | 0.19 | 0.57 | 0.98 |
Appendix A Zooniverse Results
We provide a catalog of the 157 classification results following the aggregation of votes from the Zooniverse project but prior to combining into bulk group classifications (see §3.4 for details) with this work and display the first 20 lines in Table 5. This does not include differentiation between the source of radio emission from a radio-AGN or star formation.
Appendix B Consensus Level 3 vs. Consensus Level 4 Results
@@ Appendix A Zooniverse Results
In Table 6, we present the radio-AGN statistics for consensus level 3 after radio-AGN delineation but before the combined classifications. In Table 7, we observe the consensus level 4 statistics after radio-AGN delineation and before combining classes. In Table 8, we observe the combined classification statistic version of the consensus level 4 analysis. We determine the results do not change when comparing consensus level 3 and 4. Specifically using the posterior probabilities, we find that mergers do not exhibit statistically significantly greater proportions of either C, CE, E, or Jetted sources compared to non-mergers (in other words, Pr(MNM) corresponds to a 1). Similarly to consensus level 3, we do find mergers exhibit marginally significant greater proportions of Disturbed sources in consensus level 4, where Pr(MNM) corresponds to of 2.4. We note that we observe no CE or E in the mergers. This differs from consensus level 3 as we observe no CE sources in consensus level 4, but we attribute this to be resulting from a significant sample size decrease. We also note the population proportion in non-mergers is zero. Sample size plays a significant role in consensus level 4, as it leads to more values consistent with null. While consensus level 4 has a significantly smaller sample size, it still yields similar results as consensus level 3. A change in consensus level changes our result as a higher consensus level drastically reduces our sample size. We decided consensus level 3 was the best for our study to not limit our results due to a small sample size.

### v3

@@ 6 Summary
In this work, we explored the effect of galaxy cluster mergers on the radio morphology of radio AGN. Starting from a sample of 33 galaxy clusters identified as merging (8) or nonmerging (25) by Bilton et al. (2020), we assembled a sample of 12 galaxy clusters (8 merging and 4 nonmerging) that have LoTSS 144 MHz data. Our method to determine whether the merging cluster environment has an effect on radio AGN morphology was to have 5 experts classify the sources with a radio detection using Zooniverse, according to whether a source is weak compact (WC), strong compact (SC), compact extended (CE), extended (E), FR I or FR II, FR-morphology (FRM), bent-tail (BT), or a head-tail (HT). We combined the radio morphologies classes into C (WC+SC), CE, E, jetted (FR I + FR II + FRM), and disturbed (BT + HT), ( see §4.1), and we obtain the following results:
Mergers do not exhibit a greater proportion of compact, compact extended, and jetted sources as compared to nonmergers (see §4.1).
Nonmergers contain all the extended classified sources in this study.
Mergers have a statistically significant (3) higher population proportion of Disturbed sources than nonmergers (see §4.1).
Within merging clusters there is a marginally statistically significant (2) higher proportion of AGN that are identified only as radio AGN compared to radio AGN that are also identified as optical AGN (see §4.2).
Overall, we find that the merging environment seems to have an effect on the radio morphology of radio sources in that there are more disturbed (bent and head-tails) sources in mergers. Our findings suggest the merging environment provides more opportunity for the jets to become disturbed.
@@
Within merging clusters there is a marginally statistically significant (2) higher proportion of AGN that are identified only as radio AGN compared to radio AGN that are also identified as an optical AGN (see §4.2).
Mergers do not exhibit greater proportions of radio AGNs and radio AGNs with an optical AGN detection compared to nonmergers.
This study presents significant opportunities for future work. One future direction is to follow a similar analysis to Bilton et al. (2020) and examine the spatial distribution of the classified sources in the phase-space diagram. Of particular interest is the potential trends of radio classification at certain spatial cuts within the cluster. A similar future direction would be to investigate the phase-space position of the AGNRad and AGNRad-Opt. An additional future direction is to extend this work to a larger sample of clusters and have a spectrum of cluster dynamical states. A last future direction would be the further investigation of the ”undetermined” or ”nonconsensus” morphological classifications, as there is a significant variation in the morphologies in these morphology classes.
@@ Acknowledgements
We would like to thank the anonymous referee for their useful comments that made this work better. We would also like to thank Drs. Abhijeet Borkar and Paul Martini for helpful discussions. This work was made possible by the support of the National Radio Astronomy Observatory and Green Bank Observatory, which are facilities of the US National Science Foundation operated under cooperative agreement by Associated Universities, Inc. Y.A.G. is supported by US National Science Foundation (NSF) grant AST 22-06053. This publication uses data generated via the Zooniverse.org platform, development of which is funded by generous support, including a Global Impact Award from Google, and by a grant from the Alfred P. Sloan Foundation.
LOFAR (LoTSS)
@@
Name | Cluster | Dstate | RA | DEC | Radioclass | RAGN | OAGN |
SDSS J031447.7+421321 | A0426 | M | 48.698868 | 42.222702 | WC | False | False |
SDSS J114041.6+202034 | A1367 | NM | 175.173660 | 20.342951 | WC | False | False |
SDSS J114224.4+200709 | A1367 | NM | 175.602066 | 20.119303 | CE | False | False |
SDSS J114358.1+204822 | A1367 | NM | 175.992096 | 20.806389 | WC | False | False |
SDSS J114358.9+200437 | A1367 | NM | 175.995667 | 20.077028 | CE | False | False |
SDSS J114527.6+204825 | A1367 | NM | 176.365280 | 20.807171 | CE | False | False |
SDSS J114548.8+203743 | A1367 | NM | 176.453491 | 20.628616 | WC | False | False |
SDSS J114612.1+202329 | A1367 | NM | 176.550766 | 20.391647 | SC | True | True |
SDSS J115148.4+202727 | A1367 | NM | 177.951752 | 20.457710 | WC | False | False |
SDSS J125217.7+270507 | A1656 | NM | 193.073837 | 27.085546 | WC | False | True |
SDSS J125351.5+285845 | A1656 | NM | 193.464584 | 28.979395 | WC | False | False |
SDSS J125359.1+262638 | A1656 | NM | 193.496307 | 26.444042 | WC | False | False |
SDSS J125455.1+272445 | A1656 | NM | 193.729874 | 27.412704 | WC | False | False |
SDSS J125547.8+281521 | A1656 | NM | 193.949280 | 28.256104 | WC | False | False |
SDSS J125651.1+265356 | A1656 | NM | 194.213211 | 26.898890 | WC | False | True |
SDSS J125732.8+273637 | A1656 | NM | 194.386826 | 27.610346 | WC | False | False |
SDSS J125835.1+273547 | A1656 | NM | 194.646606 | 27.596390 | SC | False | False |
SDSS J125855.9+275000 | A1656 | NM | 194.733185 | 27.833393 | CE | False | False |
SDSS J125939.1+285343 | A1656 | NM | 194.912979 | 28.895521 | WC | False | False |
SDSS J130125.2+291849 | A1656 | NM | 195.355286 | 29.313740 | CE | True | True |
Note. — Column headers are shorthand for the combined classification groups: compact (WC+SC), compact extended, extended, jetted (FRI+FRII+FRM), and disturbed (HT+BT). Rows describe the total radio AGN and proportion of the classification within nonmergers and mergers using a beta-binomial Jeffreys Bayesian prior to determine the upper and lower bound uncertainty. corresponds to the posterior probability that a specific classification is greater in mergers compared to nonmergers, discussed in Section 3.6.
Note. — Columns are radio AGN coinciding with an optical AGN determined in Bilton et al. (2020) (AGNRad-Opt) and only radio AGN (AGNRad). Rows include the total number and the proportion with beta-binomial upper and lower bound uncertainties using a Jeffreys Bayesian prior across both dynamical states. corresponds to the posterior probability that a specific classification is greater in mergers compared to nonmergers, discussed in Section 3.6.
Note. — Column names are shorthand for the classifications detailed in §3.2 weak compact (WC), strong compact (SC), compact extended (CE), extended (E), Fanaroff and Riley I+II (FRI+II), Fanaroff and Riley Morphology (FRM), bent-tail (BT), head-tail (HT), and uncertain (Unc). Rows detail the total, average (Avg), and normalized average (Norm) across mergers (M) and nonmergers (NM) for each morphology classification of consensus level of 3 after radio AGN delineation in section 3.5.
Note. — Column names are shorthand for the classifications detailed in Section §3.2 weak compact (WC), strong compact (SC), compact extended (CE), extended (E), Fanaroff and Riley I+II (FRI+II), Fanaroff and Riley Morphology (FRM), bent-tail (BT), head-tail (HT). Rows detail the total across mergers (M) and nonmergers (NM) for each morphology classification of consensus level of 4 after radio AGN delineation.
Note. — The first 20 rows of the full sample of consensus level 3 sources. The full table is provided with this article. Columns are the SDSS name of the source Bilton et al. (2020) (Name); the name of the cluster that the source is in (Cluster); the dynamical state of the host cluster, Dstate, which is either nonmerging (NM) or merging (M) as defined in Bilton et al. (2020); R.A., Decl, the radio classification assigned through the Zooniverse classification done in this work (Radio class, see §3.2); whether or not the source is classified as a radio AGN (RAGN, see §3.5); and whether or not the source is classified as an optical AGN in Bilton et al. (2020) (OAGN).
Appendix A Consensus Level 3 vs. Consensus Level 4 Results
@@
In Table 4, we present the radio AGN statistics for consensus level 3 after radio AGN delineation but before the combined classifications. In Table 5, we present the consensus level 4 statistics after radio AGN delineation and before combining classes. We determine that the results do not change when comparing consensus level 3 and 4. For example, when we compare the results for both consensus level 4 and consensus level 3 C, CE, E, or jetted sources, we find there is no difference and that both yield that mergers do not exhibit statistically significantly greater proportions of either C, CE, E, or jetted sources compared to nonmergers (in other words, Pr(MNM) corresponds to a 1). Similarly, when we compare the results for both consensus level 4 and consensus level 3 disturbed sources, we again find there is no difference in that both yield that mergers exhibit marginally significant greater proportions of disturbed sources in consensus level 4 (where Pr(MNM) is 0.982 which is equivalent to for the consensus level 4 sources). We note that in the consensus level 4 sources there are no CE in mergers. This differs from consensus level 3 sources where we observe only one CE source. We attribute the reduction of CE sources to a significant sample size decrease and is likely not physical in origin. Sample size plays a significant role in consensus level 4, as it leads to more values consistent with null. While consensus level 4 has a significantly smaller sample size, it still yields similar results as consensus level 3. While a change in consensus level does change the size of our sample, our results do not change. Thus, we decided to use the consensus level sample.
Appendix B Zooniverse Results
We provide a catalog of the 157 classification results following the aggregation of votes from the Zooniverse project but prior to combining into bulk group classifications (see §3.4 for details) with this work and display the first 20 lines in Table 6. This does not include differentiation between the source of radio emission from a radio AGN or star formation. With the information from Table 6, the reader can reproduce Table 4 and the main results from this work that are presented in Table 2 by including only those that are identified as radio AGN (see the methods of §3.5). To do this, the reader would filter Table 6 for only those with ‘RAGN’ == True.